For the following reaction, the mass of water
produced from 445 g of C57H110O6 is :
2C57H110O6(s) + 163O2(g) -----> 114CO2(g) + 110 H2O(l)
please explain and provide me the ans. as soon as possible........

Grade:11

3 Answers

Arun
4 years ago
Dear student
 
I am getting my answer as 495 gm. Please check and let me knoe if this is correct then I will share my sokution.
Mayank
4 years ago
moles of C57H110O6(s) =445/890=0.5 moles
so now we have------>
2C57H110O6(s) + 163 O2(g) -----> 114 CO2(g) + 110 H2O(l)
now moles of H2O = 110/4    (it is clear from the rxn.....)
                              =55/2
 
now mass of water =55/2 x 18=495gm  which is the ans........
hope this helps..............
 
veed vardhana
4 years ago
Yes surely the ans is 495gm.Please share your solution Arun sir as soon as possible...................................................